OpenStudy (anonymous):

(4pq)(-p^2q^3) What is the answer and why? Explain! :)

OpenStudy (mertsj):

\[4pq(-p ^{2}q ^{3})\]

OpenStudy (mertsj):

Multiply the coefficients 4 and -1. That is -4 Now multiply p and p^2. You do that by adding the exponents so p(p^2)=p^3. Now multiply q and q^3. Again, add the exponents so q(q^3) = q^4. Put it all together and you have -4p^3q^4